Give an example with justification of an element of M3 (Z) that is a unit but not the identity element.
1
Expert's answer
2020-09-25T11:00:53-0400
"Solution"

"M3(Z)" is a "3\ by\ 3" matrix over "Z"

An element "M3(Z)=(a_{ij})3\ by\ 3"

"M3(Z) = 
\begin{pmatrix}
a_{1,1} & a_{1,2} & a_{1,3} \\
a_{2,1} & a_{2,2} & a_{2,3} \\
a_{3,1} & a_{3,2} & a_{3,3} 
\end{pmatrix}"


A "3\ by\ 3" matrix over "Z" can be defined as a unit if it has all its elements as zeros except for the main diagonal, which has only ones.

"M3(Z) = 
\begin{pmatrix}
1 & 0 & 0 \\
0 & 1 & 0 \\
0 & 0 & 1 
\end{pmatrix}"

We all know that an identity matrix is the one that satisfies

"A (+, \cdot) I_{nXn}=A" Where A is any square matrix of order n

Now on additive identity of matrices, an identity is a matrix with all its elements as zeros.


"M3(Z)" is a unit but in this case it cannot be an additive identity, for example;

Let A be a matrix "\begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix}"

The additive identity

"A+M3(Z)= \begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix}+\begin{pmatrix}
1 & 0 & 0 \\
0 & 1 & 0 \\
0 & 0 & 1 
\end{pmatrix}=
\begin{pmatrix}
2 & 2 & 3 \\
2 & 4 & 4 \\
3 & 4 & 6 
\end{pmatrix} \neq
\begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix}"

But instead

"A+I_3= \begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix}+\begin{pmatrix}
0 & 0 & 0 \\
0 & 0 & 0 \\
0 & 0 & 0 
\end{pmatrix}=
\begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix} =
\begin{pmatrix}
1 & 2 & 3 \\
2 & 3 & 4 \\
3 & 4 & 5 
\end{pmatrix}"

Hence proven that an element of "M3 (Z)" that is a unit but not the identity element.